Milton needs 45 liters of 20% alcohol solution. He only has 15% alcohol solution and 30% alcohol solution. How many liters of each should he combine to make the mixture he needs?
Answer by richwmiller(17219)

a+b=45,
0.15*a+0.3*b=0.2*45
a=45-b
0.15*(45-b)+0.3*b=9
6.75-0.15b+0.3*b=9
0.15*b=2.25
b=15
a=45-b
a=30 b=15
check
0.15*30+0.3*15=0.2*45
4.5+4.5=9
9=9
ok